OpenStudy (anonymous):

Linear approximation f(x)=12-x^2; a=2 f(2.1) Got the first parts equation is y=-4x+16 Second part f(2.1) = 7.6 How do I find percent error?

OpenStudy (freckles):

sounds like you need to find the difference between the actual and approximation \[|\frac{actual-approximation}{actual} |\cdot 100 \%\]

OpenStudy (anonymous):

So what's my aproximation that's the part I don't get ?

OpenStudy (freckles):

\[f(x)=12-x^2 \approx -4x+16 \text{ for values of x near a=2}\]

OpenStudy (freckles):

-4x+16 gives approximations 12-x^2 gives actuals

OpenStudy (anonymous):

Oooo ok :) yay got it

OpenStudy (freckles):

after all the line you found is called a linear approximation

OpenStudy (freckles):

to the function given